# Importance Guided Data Augmentation for Neural-Based Code Understanding

### v1

@@ 6.2 Accuracy Evaluation
Case study 2: What factors contribute to the improved accuracy observed in code models when utilizing GenCode? Furthermore, we explore why models trained by GenCode have higher accuracy than models trained by using other code data augmentation models. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
6.3 Robustness Evaluation
@@
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of code models. First, we can see that similar to the finding in Section 6.2, not all data augmentation methods can improve the robustness of code models. Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, we can see that GenCode consistently improves the robustness of code models and achieves the best results in most cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to the SOTA data augmentation method MixCode, GenCode can train code models with 4.90% higher robustness on average. In conclusion, GenCode can produce more robust code understanding models than other methods.
@@
Case study 3: What factors contribute to the improved robustness observed in code models when utilizing GenCode? To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al. (2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 4. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods.
6.4 Ablation Study
@@
Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models in Section 4. It is still necessary to study to what extent the advantage can be brought from this selection strategy. Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values and GenCode with random selection, GenCode (Min) and GenCode (Random), respectively.
@@
Table 5 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants in both testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on all results. This indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation framework.
7 Conclusion
@@
We proposed the first generation-and-selection-based framework, GenCode, for boosting code understanding models. The basic assumption behind GenCode is that the less confident data contributes more to the model training. Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model. Our evaluation demonstrated that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively. Besides, we analyzed how the importance score affects the effectiveness of GenCode.
@@
Appendix A Datasets and Models
We evaluate GenCode on four code understanding tasks and three pre-trained code models as listed in Table 6. For all datasets, we follow the settings provided by the official projects for splitting the data into training, validation, and test sets.
  • Bug detection in source code understanding involves determining whether a piece of code contains bugs. This task can be thought of as a binary classification problem, where the target is to classify code as either bug-free or containing bugs. In our study, we utilize the Refactory dataset Hu et al. (2019), which is specifically designed for bug repair tasks. Refactory comprises 2,242 correct and 1,783 buggy Python programs. These programs were written by real-world undergraduate students to finish five different programming assignments.
  • Authorship attribution is the identification of the writer of a specific code fragment by inferring the unique characteristics of programmers from their published source code. This task serves important purposes, such as properly crediting a programmer for their contributions and aiding in the detection of code plagiarism. In our experiment, we employ a dataset sourced from Google Code Jam (GCJ) that is provided by the work Yang et al. (2022) for the authorship attribution task.
  • Clone detection is the task that focuses on determining whether two pieces of code are semantically identical. This task is instrumental in preventing the propagation of bugs and simplifying the process of software maintenance. To conduct our experiment, we adopt the widely acknowledged clone detection benchmark dataset BigCloneBench Svajlenko et al. (2014). This dataset offers an extensive and diverse set of code examples that are written by Java. Note that MixCode is not suitable for this task since its input is code pairs.
  • Problem classification is a typical task in source code understanding. It involves classifying the target functions of source code. Specifically, in a series of problems including detailed descriptions as well as their corresponding candidate source code, the trained model can correctly associate the source code with the specific problem that the code is trying to address. In our study, we leverage the recently released dataset Java250 Puri et al. (2021). Java250 is curated explicitly for Java program classification, containing a diverse set of 250 classification problems. Each problem within this dataset is paired with 300 different Java programs, each for solving this problem.
For the models, we study the powerful pre-trained code models.
  • CodeBERT Feng et al. (2020) is a bimodal model that is trained by leveraging data from various programming languages like C, C++, and Java, along with natural languages. Inspired by the principles of BERT Devlin et al. (2018), CodeBERT treats programs as sequences during pre-training.
  • GraphCodeBERT Guo et al. (2020) enhances the training data with data-flow information to incorporate the semantic-level structure of programs. This addition enriches the training process, enabling the model to generate more precise and context-aware code representations.
  • CodeT5 Wang et al. (2021) is a pre-trained encoder-decoder Transformer model that is designed to effectively capture the code semantics conveyed through developer-assigned identifiers.
Table 6: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
Dataset Language Task #Training #Test Model
Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5
Refactory Python Bug detection 3,380 423
Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5
BigCloneBench Java Clone detection 90,102 4,000
Appendix B Atomic data augmentation operators
Code Refactoring (Refactor). Code refactoring provides a range of methods aimed at altering the syntactic structure of source code while preserving its underlying semantic information. Currently, code refactoring is being increasingly employed for code data augmentation since it can be utilized to automatically create additional source code samples. Therefore, we first chose code refactoring as one of the code data augmentation methods to generate the training candidate. Furthermore, we release all 18 atomic code refactoring operations adopted by GenCode.
Synonym Replacement (SR). The original SR is applied to text data. It augments a text by randomly selecting a certain number of words and then replacing the selected words with one of their randomly chosen synonyms. Here, synonyms are machine-generated. In contrast to text data, a program is composed of a series of statements. To augment the code data, we begin by randomly selecting a set number of statements from a program. Subsequently, each word within these selected statements is randomly replaced with one of its synonyms. All operations that may alter the syntactic structure follow the rule that preserves the original relationship between the features and the label in the code.
Random Insertion (RI). For text data, RI first randomly chooses a word in a sentence and then generates a number of synonyms of the chosen word. Lastly, RI randomly selects one of the generated synonyms and inserts it into a random position of this sentence. Generally, we repeat this operation $n$ times. In code data, we first randomly choose one of the statements and generate a number of synonyms for each word in this statement. Then, we randomly choose one word and insert one of its synonyms in a random position of the chosen statement.
Random Swap (RS). RS generally randomly chooses two different words in a sentence or two different sentences in a document and then swaps their positions. This operation is usually repeated $n$ times. For source code, we first randomly select two different statements in a program and then swap their positions, also we repeat it $n$ times.
Random Deletion (RD). RD randomly selects a certain number of words in a sentence or a certain number of sentences in a document and then deletes them. For source code, we randomly delete a certain number of words in a randomly chosen statement. Similarly, this method can preserve the original relationship between the features and the label in code data.
Back-translation (BT). BT translates the original text into another language and then translates it back to the original one to generate additional text data. For code data, similarly, we choose the English-French translation model Xie et al. (2020) bidirectionally for each statement in a program.
Not that: Different from Refactor and SR, atomic data augmentation operators including RI, RS, RD, and BT can break syntax and alter the label of the code. Despite this, we found that these code augmentation operators remain beneficial in GenCode for boosting the training of code models.
Appendix C Robustness Testing of Code Model
Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data. Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness. Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data. Adversarial robustness refers to the ability of models to defend the adversarial attacks. Robustness testing of code models has been addressed recently. Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models. Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, Yang et al. (2022) proposed to change the variable name of code to force the model to make wrong predictions.
In this work, following existing works Yang et al. (2022), we employ adversarial attack methods proposed for code to evaluate the adversarial robustness of our trained model, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020).
$ASR=\frac{\#Successfully\,generated\,adversarial\,data}{\#Corrected\,classified\,test\,data}\%$ (1)
In the evaluation of robustness, ASR (in Eq.( 1)) is the metric for assessing robustness. It calculates the percentage of successfully generated adversarial examples. Adversarial attacks are performed exclusively on correctly predicted test samples. A higher ASR suggests that an attack method is more effective, indicating lower robustness of the victim model.

### v2

@@ Introduction
Visualizing principal component analysis. Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak 1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure. 3. The results from Table 4 confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
Robustness Evaluation
@@
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of adversarial attack methods against code models. First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models. This finding is consistent with the existing works (Dong et al. 2023b). Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding (Bielik and Vechev 2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
Statistical analysis. We also conduct statistical tests using T-Test and Wilcoxon signed-rank test. When comparing existing data augmentation methods, GenCode consistently exhibits a p-value $<$ 0.05 across all pre-trained models. In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
@@
Maximum probability analysis. To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data (Wu et al. 2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 5. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods. As a result, GenCode trained code models are harder to attack.
Ablation Study
@@
Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models, it is still necessary to study to what extent the advantage can be brought from this selection strategy. Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values - GenCode (Min) and random selection - GenCode (Random), respectively.
@@
Table 6 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants under all testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results. This phenomenon indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation method.
Discussion
Even though the accuracy of our studied code models already achieved relatively high accuracy in these tasks, i.e., most of them have more than 90% accuracy, their robustness is still low. Regarding the large language models (LLMs), we have checked the literature and found that the clone detection task is still challenging for LLMs, e.g., ChatGPT has only 0.784 Precision for Java-Java pair detection (Khajezade et al. 2024). For other program understanding tasks that have a large number of classes, e.g., the Java250 dataset used in our experiments has 250 classes. It is also hard to use open-source LLMs such as ChatGPT to predict the label since we need to add many label guidance in the prompt.
Conclusion
@@
We propose the first generation-and-selection-based framework, GenCode, to boost code understanding models. The basic assumption behind GenCode is that the less confident data contributes more to the model training. Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model. Our evaluation demonstrates that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively. Besides, we analyze how the importance score affects the effectiveness of GenCode.